a particle moves from a point  r1 =(2m)i+(3m)j to another point r2 = (3m)i +(2m)j during which a certain force F=(5N)i+(5N)j  acts on it .find the work done by the force on the particle during the displacement.

Grade:12

7 Answers

KS
11 years ago

Work done depends upon path taken,

There are two possible paths

1) (2)i

mohit yadav
11 years ago

the work done is 0

harihara kumar
11 years ago

w=f.(r2-r1)=0

Akash Kumar Dutta
11 years ago

displacement=S=r2-r1=(m)i - (m)j
WD=F.S= (5m)i - (5m)j
|WD|= root(50m^2)=5mroot2 ANS

Chetan Mandayam Nayakar
11 years ago

answer=(5Ni+5Nj)·(r2-r1)

vivek kumar
10 years ago

work done is zero.net displacement =i-j.

and work=f.ds net work=5-5=0.

SOURAV MISHRA
10 years ago

the displacement of the particle is r2 - r1 = i - j.

therefore the work done is F.(r2 - r1) = 0 J.

so the work done is zero.